A 2 kg block is held in equilibrium on a incline of angle 60° by horizontal force F aplied in the direction. if the coefficient between block and incline is 0.3. determine the minimum value of horizontal force F.
1
Expert's answer
2020-10-21T09:53:22-0400

Draw the free body diagram:



Write the equilibrium of forces by axes:


"f=\mu N\\
x:f+F\text{ cos}\alpha-mg\text{ sin}\alpha=0,\\
y: N-mg\text{ cos}\alpha-F\text{ sin}\alpha=0."

From the equations, we have


"F=\frac{mg(\text{sin}\alpha-\mu\text{cos}\alpha)}{\mu\text{sin}\alpha+\text{cos}\alpha}=18.5\text{ N}."